# Pimba: A Processing-in-Memory Acceleration for Post-Transformer Large Language Model Serving

### v2

@@ Abstract.
Transformers are the driving force behind today’s Large Language Models (LLMs), serving as the foundation for their performance and versatility. Yet, their compute and memory costs grow with sequence length, posing scalability challenges for long-context inferencing. In response, the algorithm community is exploring alternative architectures—such as state space models (SSMs) (e.g., Mamba-2), linear attention, and recurrent neural networks (RNNs)—which we refer to as post-transformers. This shift presents a key challenge: building a serving system that efficiently supports both transformer and post-transformer LLMs within a unified framework.
To address this challenge, we analyze the performance characteristics of transformer and post-transformer LLMs. Despite their algorithmic differences, both are fundamentally limited by memory bandwidth under batched inference—due to attention in transformers and state updates in post-transformers. Inspired by this finding, we propose Pimba, an accelerator solution that aims to address the memory bottleneck by jointly leveraging (1) Processing-in-Memory (PIM) paradigm and (2) LLM quantization. Further analyses suggest two additional insights: (1) state update operations, unlike attention, incur high hardware cost, making per-bank PIM acceleration inefficient, and (2) different low-precision arithmetic methods offer varying accuracy-area tradeoffs, while we identify Microsoft’s MX as the Pareto-optimal choice. Building on these insights, we design Pimba as an array of State-update Processing Units (SPUs), each shared between two banks to enable interleaved access to PIM. Each SPU includes a State-update Processing Engine (SPE) that comprises element-wise multipliers and adders using MX-based quantized arithmetic, enabling efficient execution of state update and attention operations. Our evaluation shows that, compared to LLM-optimized GPU and GPU+PIM systems, Pimba achieves up to 4.1 and 2.1 higher token generation throughput, respectively.
@@ 1. Introduction
While transformers offer impressive capabilities and continue to dominate LLMs, their enormous resource demands are a significant concern for LLM serving providers. Transformer-based LLMs scale quadratically in compute and linearly in memory footprint with sequence length, while emerging applications–such as test time scaling (DeepSeek-AI et al., 2025; Snell et al., 2025; o1, [n. d.]; Muennighoff et al., 2025), retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022; Lan et al., 2023; Schick et al., 2023), and multimodal input fusion (gpt, [n. d.]; lla, [n. d.])–are driving demand for longer sequence lengths, recently reaching up to 2 billion tokens in state-of-the-art models (Yiran Ding and Li Lyna Zhang and Chengruidong Zhang and Yuanyuan Xu and Ning Shang and Jiahang Xu and Fan Yang and Mao Yang, 2024). Moreover, batched inferencing exacerbates these resource demands, forcing hyperscalers to invest billions of dollars in equipping their data centers with hundreds of thousands of costly GPUs, each priced at $30,000 or more (PCMag, [n. d.]).
@@
Recently, the algorithm community has actively explored alternative approaches, including state space models (SSMs), linear attention mechanisms, and recurrent neural networks (RNNs). In this paper, we henceforth refer to LLMs employing these alternative algorithmic techniques as “post-transformer” LLMs. We argue that post-transformer LLMs have the potential to serve as a promising complement to transformer-based LLMs, providing comparable algorithmic capabilities with significantly lower and constant resource demands (Gu and Dao, 2024; Dao and Gu, 2024; Gu et al., 2022; Fu et al., 2023; Poli et al., 2023; Gu et al., 2024, 2020, 2021; Katharopoulos et al., 2020; Yang et al., 2024b; Zhang et al., 2024; Sun et al., 2023; Peng et al., 2023, 2024; Beck et al., 2024; Qin et al., 2024). Figure 1(a) presents an empirical evidence that supports our argument. The figure compares the memory usage, throughput, and accuracy of a transformer-based LLM with a post-transformer LLM, Mamba-2, both having a model size of 2.7 billion parameters111Detailed experimental methodology is presented in Section 6.1. The results show that Mamba-2 requires 2.3 less memory capacity, delivering 2.6 higher throughput than the transformer counterpart, while achieving 4.5% higher accuracy. Despite this considerable potential, the architecture and system community have a limited understanding of the implications of these algorithms, causing LLM service providers to hesitate in adopting them for their serving systems.
@@
To this end, this paper sets out to bridge this gap through a comprehensive workload analysis and performance characterization, leveraging the findings and insights to devise an acceleration solution. The first crucial finding is that many post-transformer LLMs share a common algorithmic operation, state update, which propagates and evolves contextual information across tokens. This commonality offers a promising opportunity for architectural generalization and acceleration. We also discovered that similar to the attention operation in transformer-based LLMs, this state update operation becomes the performance bottleneck due to its low arithmetic intensity. Figure 1(b) reports the roofline analysis results that the arithmetic intensity of state update operation is 4 larger than that of attention, while it is still significantly bandwidth bound.
@@
(Principle 1): Maximizing hardware resource sharing for area efficiency: Existing LLM-targeted PIM acceleration methods (Lee et al., 2021; He et al., 2020; Heo et al., 2024; Park et al., 2024a; Seo et al., 2024) focus on supporting matrix-vector multiplication (i.e., GEMV) since attention operation is consisted of full of GEMVs. However, this approach is unsuitable for post-transformer algorithms since implementing the state update operation in hardware incurs significantly larger area costs due to the variety of primitives in state update operation, such as element-wise multiplication, element-wise addition, and vector dot products. Thus, in designing Pimba, we aim to exploit the hardware resource sharing for maximizing area efficiency.
(Principle 2): Achieving both accuracy and area-efficiency from low-precision arithmetic: While quantizing the state in post-transformers can reduce computation cost and memory footprint, it also affects area efficiency. We also discovered that, due to the state “update” mechanism, conventional numerical formats cause severe accuracy degradation, rendering them impractical for post-transformers. We carefully explore the accuracy-area tradeoffs and observe that different low-precision arithmetic approaches exhibit different characteristics. We thoroughly perform an empirical study to understand the differences and aim to employ a Pareto-optimal quantization technique for our solution.
@@
End-to-end Pimba system design. We construct the Pimba system by jointly leveraging Pimba accelerators with GPUs, offloading state update and attention operations to PIM, while delegating other tasks to GPUs. Pimba includes custom DRAM commands and command scheduling techniques to manage state pre-charging and subsequent generative computations. Our PIM accelerator and its interface use a system architecture similar to existing PIM-based LLM serving systems (Heo et al., 2024; Park et al., 2024a; Seo et al., 2024; Park et al., 2024b), allowing Pimba to serve as a “drop-in replacement” in transformer-serving systems adapted to support post-transformer LLMs as well.
@@ 2.1. Transformer-based LLMs and their Limitations
Memory usage. The KV cache grows linearly with the sequence length. Despite prior work on enhancing memory efficiency (Hooper et al., 2024; Yang et al., 2024a; Dong et al., 2024a; Kang et al., 2024; Kwon et al., 2023), the fundamental property of the algorithm is unchanged. This ends up consuming significant amounts of GPU memory and imposing limits on the sequence length or the batch size.
@@ 2.2. Post-Transformer LLMs
Combining with attention. Although aforementioned architectures demonstrate strong performance, they often fall short in in-context learning, particularly in recalling previous tokens (Waleffe et al., 2024). This limitation has motivated a body of work (NVIDIA, [n. d.]; Zyphra, [n. d.]; Dong et al., 2024b; Ren et al., 2025; Waleffe et al., 2024) exploring hybrid models that combine the efficiency of alternative architectures and the expressiveness of attention. Notably, Nemotron-H (NVIDIA, [n. d.]) and Zamba2 (Zyphra, [n. d.]) integrate attention layers with Mamba-2 architectures to leverage the complementary strengths of both approaches. By sparsely inserting attention layers, for example, one attention layer per six Mamba-2 layers in Zamba2 (Zyphra, [n. d.]), these models effectively restore the in-context learning capability of standard Transformers, while maintaining the computational efficiency.
@@ Acknowledgements.
We thank the anonymous reviewers for their comments and feedback. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) (No.RS-2024-00396013), IITP under the Graduate School of Artificial Intelligence Semiconductor (IITP-2025-RS-2023-00256472), grant funded by the Korea government (MSIT). This work was also supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by ICT R&D program of MSIT/IITP (No.RS-2025-02305453, Development of distributed inference and model optimization technology for heterogeneous Al semiconductors). The EDA tool was supported by the IC Design Education Center (IDEC), Korea.References
@@
- Dong et al. (2024a) Shichen Dong, Wen Cheng, Jiayu Qin, and Wei Wang. 2024a. QAQ: Quality Adaptive Quantization for LLM KV Cache. https://arxiv.org/abs/2403.04643
- Dong et al. (2024b) Xin Dong, Yonggan Fu, Shizhe Diao, Wonmin Byeon, Zijia Chen, Ameya Sunil Mahabaleshwarkar, Shih-Yang Liu, Matthijs Van Keirsbilck, Min-Hung Chen, Yoshi Suhara, Yingyan Lin, Jan Kautz, and Pavlo Molchanov. 2024b. Hymba: A Hybrid-head Architecture for Small Language Models. https://arxiv.org/abs/2411.13676
@@
- Gu and Dao (2024) Albert Gu and Tri Dao. 2024. Mamba: Linear-Time Sequence Modeling with Selective State Spaces. https://arxiv.org/abs/2312.00752
@@
- Muennighoff et al. (2025) Niklas Muennighoff, Zitong Yang, Weijia Shi, Xiang Lisa Li, Li Fei-Fei, Hannaneh Hajishirzi, Luke Zettlemoyer, Percy Liang, Emmanuel Candès, and Tatsunori Hashimoto. 2025. s1: Simple test-time scaling. https://arxiv.org/abs/2501.19393
@@
- Tianqi et al. (2025) Chen Tianqi, Yuanteng Chen, Weixiang Xu, Zeyu Zhu, Peisong Wang, and Jian Cheng. 2025. Q-Mamba: Towards more efficient Mamba models via Post-Training Quantization. https://openreview.net/forum?id=AY1S52vr0a

### v3

@@ Abstract.
Transformers are the driving force behind today’s Large Language Models (LLMs), serving as the foundation for their performance and versatility. Yet, their compute and memory costs grow with sequence length, posing scalability challenges for long-context inferencing. In response, the algorithm community is exploring alternative architectures—such as state space models (SSMs) (e.g., Mamba-2), linear attention, and recurrent neural networks (RNNs)—which we refer to as post-transformers. This shift presents a key challenge: building a serving system that efficiently supports not only emerging post-transformer LLMs but also existing transformer models within a unified framework.
To address this challenge, we analyze the performance characteristics of transformer and post-transformer LLMs. Despite their algorithmic differences, both are largely bounded by memory bandwidth under batched inference—due to attention in transformers and state updates in post-transformers. Inspired by this finding, we propose Pimba, an accelerator solution that aims to address the memory bottleneck by jointly leveraging (1) Processing-in-Memory (PIM) paradigm and (2) LLM quantization. Further analyses suggest two additional insights: (1) state update operations, unlike attention, incur high hardware cost, making per-bank PIM acceleration inefficient, and (2) different low-precision arithmetic methods offer varying accuracy-area tradeoffs, while we identify Microsoft’s MX as a Pareto-optimal choice. Building on these insights, we design the architecture of Pimba as an array of State-update Processing Units (SPUs), each shared between two banks to enable interleaved access. Each SPU includes a State-update Processing Engine (SPE) that comprises element-wise multipliers and adders using MX-based quantized arithmetic, enabling efficient execution of state update and attention operations. Our evaluation shows that, compared to LLM-optimized GPU and GPU+PIM systems, Pimba achieves up to 4.1 and 2.1 higher generation throughput, respectively.
@@ 1. Introduction
While transformers offer remarkably versatile capabilities and continue to dominate LLMs, their enormous resource demands are a significant concern for LLM providers. Transformer-based LLMs scale quadratically in compute and linearly in memory footprint with sequence length, while emerging applications–such as test time scaling (DeepSeek-AI et al., 2025; Snell et al., 2025; o1, [n. d.]; Muennighoff et al., 2025), retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022; Lan et al., 2023; Schick et al., 2023), and multimodal input fusion (gpt, [n. d.]; lla, [n. d.])–are driving demand for longer sequence lengths, recently reaching up to 2 billion tokens in industry-leading models (Yiran Ding and Li Lyna Zhang and Chengruidong Zhang and Yuanyuan Xu and Ning Shang and Jiahang Xu and Fan Yang and Mao Yang, 2024). Moreover, batched inferencing exacerbates these resource demands, forcing hyperscalers to invest billions of dollars in equipping their data centers with hundreds of thousands of costly GPUs, each priced at $30,000 or more (PCMag, [n. d.]).
@@
Recently, the algorithm community has actively explored alternative approaches, including state space models (SSMs), linear attention mechanisms, and recurrent neural networks (RNNs). In this paper, we henceforth refer to LLMs employing these alternative algorithmic techniques as “post-transformer” LLMs. We argue that post-transformer LLMs have the potential to serve as a promising complement to transformer-based LLMs, providing comparable algorithmic capabilities with significantly lower and constant resource demands (Gu and Dao, 2024; Dao and Gu, 2024; Gu et al., 2022; Fu et al., 2023; Poli et al., 2023; Gu et al., 2024, 2020, 2021; Katharopoulos et al., 2020; Yang et al., 2024b; Zhang et al., 2024; Sun et al., 2023; Peng et al., 2023, 2024; Beck et al., 2024; Qin et al., 2024). Figure 1(a) presents an empirical evidence that supports our argument. The figure compares the memory usage, throughput, and accuracy of a transformer-based LLM with a post-transformer LLM, Mamba-2, both having a model size of 2.7 billion parameters111Detailed experimental methodology is presented in Section 6.1. The results show that Mamba-2 requires 2.3 less memory capacity, delivering 2.6 higher throughput than the transformer counterpart, while achieving 4.5% higher accuracy. Despite this considerable potential, the architecture and system community have a limited understanding of the implications of these algorithms, causing LLM providers to hesitate in adopting them for their serving systems.
@@
To this end, this paper sets out to bridge this gap through a comprehensive workload analysis and performance characterization, and to devise a solution that leverages the resulting insights. The first of these insights is that many post-transformer LLMs share a common algorithmic operation, state update, which propagates and evolves contextual information across tokens. This commonality offers a promising opportunity for architectural generalization and acceleration. We also discovered that similar to the attention operation in transformer-based LLMs, this state update operation becomes the performance bottleneck due to its low arithmetic intensity. Figure 1(b) reports the roofline analysis results that the arithmetic intensity of state update operation is 4 larger than that of attention, while it is still significantly bandwidth-bound.
@@
(Principle 1): Maximizing hardware resource sharing for area efficiency: Existing LLM-targeted PIM acceleration methods (Lee et al., 2021; He et al., 2020; Heo et al., 2024; Park et al., 2024a; Seo et al., 2024) focus on supporting matrix-vector multiplication (i.e., GEMV) since attention operation consists of a full of GEMVs. However, this approach is unsuitable for post-transformer algorithms since implementing the state update operation in hardware incurs significantly larger area costs due to the variety of primitives in state update operation, such as element-wise multiplication, element-wise addition, and vector dot products. Thus, in designing Pimba, we aim to exploit the hardware resource sharing for maximizing area efficiency.
(Principle 2): Achieving both accuracy and area-efficiency from low-precision arithmetic: While quantizing the state in post-transformers can reduce computation cost and memory footprint, it also affects area efficiency. We also discover that, due to the state “update” mechanism, conventional numerical formats cause severe accuracy degradation, rendering them impractical for post-transformers. We carefully explore the accuracy-area tradeoffs and observe that different low-precision arithmetic approaches exhibit different characteristics. We thoroughly perform an empirical study to understand the differences and aim to employ a Pareto-optimal quantization technique for our solution.
@@
End-to-end Pimba system design. We construct the Pimba system by jointly leveraging Pimba accelerators with GPUs, offloading state update and attention operations to PIM, while delegating other tasks to GPUs. Pimba includes custom DRAM commands and command scheduling techniques to manage state pre-charging and subsequent generative computations. Our PIM accelerator and its interface use a system architecture similar to the existing PIM-based LLM serving systems (Heo et al., 2024; Park et al., 2024a; Seo et al., 2024; Park et al., 2024b), allowing Pimba to serve as a “drop-in replacement” in transformer-serving systems adapted to support post-transformer LLMs as well.
@@ 2.1. Transformer-based LLMs and their Limitations
Memory usage. The KV cache grows linearly with the sequence length. Despite prior work on enhancing memory efficiency (Hooper et al., 2024; Yang et al., 2024a; Dong et al., 2024; Kang et al., 2024; Kwon et al., 2023), the fundamental property of the algorithm is unchanged. This ends up consuming significant amounts of GPU memory and imposing limits on the sequence length or the batch size.
@@ 2.2. Post-Transformer LLMs
Combining with attention. Although aforementioned architectures demonstrate strong performance, they often fall short in in-context learning, particularly in recalling previous tokens (Waleffe et al., 2024). This limitation has motivated a body of work (NVIDIA, [n. d.]; Zyphra, [n. d.]; Dong et al., 2025; Ren et al., 2025; Waleffe et al., 2024) exploring hybrid models that combine the efficiency of alternative architectures and the expressiveness of attention. Notably, Nemotron-H (NVIDIA, [n. d.]) and Zamba2 (Zyphra, [n. d.]) integrate attention layers with Mamba-2 architectures to leverage the complementary strengths of both approaches. By sparsely inserting attention layers, for example, one attention layer per six Mamba-2 layers in Zamba2 (Zyphra, [n. d.]), these models effectively restore the in-context learning capability of standard Transformers, while maintaining the computational efficiency.
@@ Acknowledgements.
We thank the anonymous reviewers for their comments and feedback. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) (No.RS-2024-00396013, No.2022-0-01037), IITP under the Graduate School of Artificial Intelligence Semiconductor (IITP-2025-RS-2023-00256472), grant funded by the Korea government (MSIT). This work was also supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by ICT R&D program of MSIT/IITP (No.RS-2025-02305453, Development of distributed inference and model optimization technology for heterogeneous Al semiconductors). The EDA tool was supported by the IC Design Education Center (IDEC), Korea.References
@@
- Astral ([n. d.]) Astral. [n. d.]. An extremely fast Python package and project manager, written in Rust. https://docs.astral.sh/uv/.
@@
- Dong et al. (2024) Shichen Dong, Wen Cheng, Jiayu Qin, and Wei Wang. 2024. QAQ: Quality Adaptive Quantization for LLM KV Cache. https://arxiv.org/abs/2403.04643
- Dong et al. (2025) Xin Dong, Yonggan Fu, Shizhe Diao, Wonmin Byeon, ZIJIA CHEN, Ameya Sunil Mahabaleshwarkar, Shih-Yang Liu, Matthijs Van keirsbilck, Min-Hung Chen, Yoshi Suhara, Yingyan Celine Lin, Jan Kautz, and Pavlo Molchanov. 2025. Hymba: A Hybrid-head Architecture for Small Language Models. In ICLR.
@@
- Gu and Dao (2024) Albert Gu and Tri Dao. 2024. Mamba: Linear-Time Sequence Modeling with Selective State Spaces. In COLM.
@@
- Muennighoff et al. (2025) Niklas Muennighoff, Zitong Yang, Weijia Shi, Xiang Lisa Li, Li Fei-Fei, Hannaneh Hajishirzi, Luke Zettlemoyer, Percy Liang, Emmanuel Candes, and Tatsunori Hashimoto. 2025. s1: Simple test-time scaling. In Workshop on Reasoning and Planning for Large Language Models.
@@
- Tianqi et al. (2025) Chen Tianqi, Yuanteng Chen, Peisong Wang, Weixiang Xu, Zeyu Zhu, and Jian Cheng. 2025. Q-Mamba: Towards more efficient Mamba models via post-training quantization. In ACL.
@@
Appendix A Artifact Appendix
A.1. Abstract
In this work, we present Pimba, a processing-in-memory acceleration solution for post-transformer large language models. Our artifact includes a full-system simulator for Pimba as well as accuracy evaluation code. To ensure reproducibility and streamline the experimental workflow, we made two key engineering efforts: (1) The entire codebase relies solely on the modern project manager uv, allowing evaluators to install exactly the same versions of dependencies and build tools with a one-line command; and (2) we provide only two scripts–one to run all experiments and another to generate all figures from the results–to reduce the burden on evaluators. These two engineering efforts collectively enable a streamlined experimental workflow and maximize reproducibility.
Beyond the artifact evaluation, we also document the challenges we encountered while preparing the code and a breif API reference to assist those who wish to extend our codebase. We believe these resources will help those preparing artifact evaluations and aid future research efforts.
A.2. Artifact check-list
Program: uv
Compilation: gcc
Model: Our code automatically downloads all required model weights from Hugging Face: RetNet, GLA, HGRN2, Mamba2, Zamba2, OPT, LLaMA
Data sets: Our code automatically downloads all required datasets from Hugging Face: Wikitext2, Piqa, Lambada, Hellaswag, Arc-easy, Arc-challenge, Winogrande
Run-time environment: CUDA, x86_64
Hardware: An NVIDIA GPU based on the Ampere architecture or newer, equipped with at least 24GB of memory
Metrics: Perplexity, Accuracy, Throughput
Output: PDF files for the figures, a CSV file for the table
How much disk space required (approximately)?: The installation occupies about 100GB in the user’s ”$HOME” directory, primarily due to models and datasets stored under ”$HOME/.cache/huggingface”.
How much time is needed to prepare workflow (approximately)?: 10 minutes
How much time is needed to complete experiments (approximately)?: 15 hours
Publicly available?: https://github.com/casys-kaist/pimba
Archived (provide DOI)?: https://doi.org/10.5281/zenodo.16946084
A.3. Description
How to access. Clone our public repository from github:
$ git clone https://github.com/casys-kaist/pimba
Hardware dependencies. Evaluators require an NVIDIA GPU based on the Ampere architecture or newer, with at least 24GB of memory (e.g., RTX 3090, RTX 4090, RTX 5090, A6000, A100, etc.).
Software dependencies. Evaluators only need gcc and uv; all other required python packages and build tools are installed through uv. uv itself can be installed using the system package manager or via a one-line command provided in the official documentation (Astral, [n. d.]).
Models and data sets. During the experiments, our code automatically downloads all required models and datasets.
A.4. Installation
With the following commands, uv automatically downloads the required dependencies and build tools (e.g., cmake, ninja) in exactly the same versions we used and compiles our project.
$ uv sync $ uv run cmake --preset release $ uv run cmake --build build
A.5. Experiment workflow
Evaluators can run all experiments with a single command:
$ uv run python scripts/run.py
This command generates two files, ”accuracy_result.yaml” and ”performance_result.yaml”, under the ”res/” directory, which are then used to reproduce the figures and the table.
A.6. Evaluation and expected results
Using the result files, evaluators can simply reproduce the figures and the table in the paper by executing the following command:
$ uv run python scripts/draw.py
This process generates PDF files for the figures and a CSV file for the table in the ”summary/” directory. Note that the PDF files do not include axis labels or legends, as we add them using external drawing tools. However, all scales and colors precisely match those in the paper, ensuring no issues arise when comparing values.
We also observed that accuracy may slightly vary with the GPU platform and CUDA version used for evaluation, especially in Table 2. However, the key trend we aim to convey in this paper remains consistent: quantization with MX8 has minimal impact on accuracy.